OpenStudy (anonymous):

factorizen 4(5x+1/3)^2+1=101

OpenStudy (anonymous):

subtract 1 from both sides divide both sides by 4 take the square root, don't forget the \(\pm\) then solve for \(x\)

OpenStudy (anonymous):

not sure what "factorizen" means, but it should be "solve for \(x\)"

OpenStudy (tkhunny):

4(5x+1/3)^2+1=101 Subtract 101 4(5x+1/3)^2 - 100 = 0 Factor out 4 4[(5x+1/3)^2 - 25] = 0 Recognize the Difference of Squares 4[5x+1/3 + 5][5x+1/3 - 5] = 0 Minor adjustments 4[5x + 16/3][5x - 14/3] = 0 Factor out the 5 if you wish. 20[x + 16/15][x - 14/15] = 0

OpenStudy (anonymous):

that will work as well, but i think it is more complicated

OpenStudy (anonymous):

you judge what you like for yourself \[4(5x+\frac{1}{3})^2+1=101 \\4(x+\frac{1}{3})^2=100\\ (5x+\frac{1}{3})^2=25\\ 5x+\frac{1}{3}=5,5x+\frac{1}{3}=-5\] etc